OpenStudy (anonymous):

which equation represents the axis of symmetry of the graph pf the equation y=x^2-4x-12

OpenStudy (aravindg):

Convert it into the standard form of parabola \[(x-h)^2=4a(y-k)\]

OpenStudy (aravindg):

Then put 2nd order term=0 to get axis of symmetry.

Parth (parthkohli):



Parth (parthkohli):

The axis of symmetry is a vertical line passing through the vertex. The x-coordinate of the vertex is \(\frac{-b}{2a}\). So your axis of symmetry would be \(x=\frac{-b}{2a}\).

Parth (parthkohli):

That's simpler, right?

Parth (parthkohli):

The graph that I drew there does not represent the parabola y = x^2 - 4x - 12.